# Improving Video Generation with Human Feedback

### v1

@@ A.1 Relation beween Velocity and Noise
Let $X_{0}\sim q$ be a real data sample drawn from the true data distribution and $X_{1}\sim p$ be a noise sample, where $X_{0},X_{1}\in\mathbb{R}^{d}$. Define ${\bm{v}}_{t}({\bm{x}}_{t}\mid X_{0},X_{1})$ to be the conditional velocity field specified by a Rectified Flow (Liu et al., 2022), and let ${\bm{v}}_{t}^{\text{pred}}({\bm{x}}_{t})$ be the predicted marginal velocity field. Then the L2 error of the noise prediction is related to the L2 error of the velocity field prediction by
@@
=\;= $\mathbb{E}\bigl{[}\dot{\psi}(X_{t}\mid X_{0},X_{1})\,\bigm{|}\,X_{t}={\bm{x}}_{t}\bigr{]}$
=\;= $\mathbb{E}\bigl{[}X_{1}-X_{0}\,\bigm{|}\,X_{t}={\bm{x}}_{t}\bigr{]}$
[…]
@@
A.2 Reward as Classifier Guidance
We begin by citing a lemma from the Guided Flows paper (Zheng et al., 2023)
Lemma A.2.
@@ A.2 Reward as Classifier Guidance
Seed Appendix A.61 of the Guided Flows paper (Zheng et al., 2023) for detailed derivations.
@@
$\tilde{v}_{t}({\bm{x}}_{t}|{\bm{y}})$ $={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[a_{t}{\bm{x}}_{t}+b_{t}\nabla r({\bm{x}}_{t},{\bm{y}})-{\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})]$ (23)
$={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[a_{t}{\bm{x}}_{t}+b_{t}\nabla\log\exp(r({\bm{x}}_{t},{\bm{y}}))-{\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})]$ (24)
$=(1-w){\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[a_{t}{\bm{x}}_{t}+b_{t}\nabla\log\exp(r({\bm{x}}_{t},{\bm{y}}))]$ (25)
$=a_{t}{\bm{x}}_{t}+b_{t}\nabla\left[(1-w)\log p_{t}({\bm{x}}_{t}|{\bm{y}})+w\log\exp(r({\bm{x}}_{t},{\bm{y}}))\right]$ (26)
@@
We note, however, that this analysis shows that Reward Guided Flows are guaranteed to sample from $\tilde{q}(\cdot|{\bm{y}})$ at time $t=1$ if the probability path $\tilde{p}_{t}(\cdot|{\bm{y}})$ is close to the marginal probability path $\int p_{t}(\cdot|{\bm{x}}_{1})\tilde{q}({\bm{x}}_{1}|{\bm{y}})dx_{1}$, but it is not clear to what extent this assumption holds in practice. This also mens that $\tilde{p_{t}}({\bm{x}}_{t}|{\bm{y}})$ is also a marginal gaussian path defined by $p_{t}({\bm{x}}|{\bm{x}}_{1})=\mathcal{N}({\bm{x}}|\alpha_{t}{\bm{x}}_{1},\sigma_{t}^{2}I)$.
@@
$\tilde{v}_{t}({\bm{x}}_{t}|{\bm{y}})$ $={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+wb_{t}\nabla r({\bm{x}}_{t},{\bm{y}})$ (28)
$={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+wb_{t}\nabla\log\exp(r({\bm{x}}_{t},{\bm{y}}))$ (29)
$=a_{t}{\bm{x}}_{t}+b_{t}\nabla\left[\log p_{t}({\bm{x}}_{t}|{\bm{y}})+w\log\exp(r({\bm{x}}_{t},{\bm{y}}))\right]$ (30)
@@
Rectified Flow (Liu et al., 2022) is also a Gaussian path defined by
@@
We use Eq. 24 or Eq. 29 to guide inference through reward model.
Appendix B Input Template for Reward Model
Full Input Template [VIDEO] You are tasked with evaluating a generated video based on three distinct criteria: Visual Quality, Motion Quality, and Text Alignment. Please provide a rating from 0 to 10 for each of the three categories, with 0 being the worst and 10 being the best. Each evaluation should be independent of the others.
**Visual Quality:**
Evaluate the overall visual quality of the video, with a focus on static factors. The following sub-dimensions should be considered:
- **Reasonableness:** The video should not contain any significant biological or logical errors, such as abnormal body structures or nonsensical environmental setups.
- **Clarity:** Evaluate the sharpness and visibility of the video. The image should be clear and easy to interpret, with no blurring or indistinct areas.
- **Detail Richness:** Consider the level of detail in textures, materials, lighting, and other visual elements (e.g., hair, clothing, shadows).
- **Aesthetic and Creativity:** Assess the artistic aspects of the video, including the color scheme, composition, atmosphere, depth of field, and the overall creative appeal. The scene should convey a sense of harmony and balance.
- **Safety:** The video should not contain harmful or inappropriate content, such as political, violent, or adult material. If such content is present, the image quality and satisfaction score should be the lowest possible.
Please provide the ratings of Visual Quality: <|VQ_reward|>
END
**Motion Quality:**
Assess the dynamic aspects of the video, with a focus on dynamic factors. Consider the following sub-dimensions:
- **Stability:** Evaluate the continuity and stability between frames. There should be no sudden, unnatural jumps, and the video should maintain stable attributes (e.g., no fluctuating colors, textures, or missing body parts).
- **Naturalness:** The movement should align with physical laws and be realistic. For example, clothing should flow naturally with motion, and facial expressions should change appropriately (e.g., blinking, mouth movements).
- **Aesthetic Quality:** The movement should be smooth and fluid. The transitions between different motions or camera angles should be seamless, and the overall dynamic feel should be visually pleasing.
- **Fusion:** Ensure that elements in motion (e.g., edges of the subject, hair, clothing) blend naturally with the background, without obvious artifacts or the feeling of cut-and-paste effects.
- **Clarity of Motion:** The video should be clear and smooth in motion. Pay attention to any areas where the video might have blurry or unsteady sections that hinder visual continuity.
- **Amplitude:** If the video is largely static or has little movement, assign a low score for motion quality.
Please provide the ratings of Motion Quality: <|MQ_reward|>
END
**Text Alignment:**
Assess how well the video matches the textual prompt across the following sub-dimensions:
- **Subject Relevance** Evaluate how accurately the subject(s) in the video (e.g., person, animal, object) align with the textual description. The subject should match the description in terms of number, appearance, and behavior.
- **Motion Relevance:** Evaluate if the dynamic actions (e.g., gestures, posture, facial expressions like talking or blinking) align with the described prompt. The motion should match the prompt in terms of type, scale, and direction.
- **Environment Relevance:** Assess whether the background and scene fit the prompt. This includes checking if real-world locations or scenes are accurately represented, though some stylistic adaptation is acceptable.
- **Style Relevance:** If the prompt specifies a particular artistic or stylistic style, evaluate how well the video adheres to this style.
- **Camera Movement Relevance:** Check if the camera movements (e.g., following the subject, focus shifts) are consistent with the expected behavior from the prompt.
Textual prompt - [PROMPT]
Please provide the ratings of Text Alignment: <|TA_reward|>
END
Appendix C Details of Human Annotation
@@
Appendix D Details of Reward Model Evaluation
D.1 Evaluation Benchmarks
We use two benchmarks to evaluate the performance of our reward model: GenAI-Bench (Jiang et al., 2024b) and VideoGen-RewardBench. GenAI-Bench is employed to assess the accuracy of the reward model in evaluating pre-SOTA-era T2V models, while VideoGen-RewardBench is used to evaluate its performance on modern T2V models. In this subsection, we provide a detailed description of the two benchmarks, highlighting their key parameters and differences in Fig 8 and Tab. 9. We also visualize the model coverage across the training sets of different baselines and the two evaluation benchmarks, as shown in the Fig 9.
@@ D.1 Evaluation Benchmarks
VideoGen-Eval (Zeng et al., 2024) has open-sourced a dataset containing videos generated by 9 closed-source and 3 open-source models, designed to qualitatively visualize performance differences across models. Due to its high-quality data, broad coverage of the latest advancements in T2V models, and third-party nature, we leverage VideoGen-Eval to create a fair benchmark, VideoGen-RewardBench, for evaluating reward models’ performance on modern T2V models. We manually construct 26.5k video pairs and hire annotators to assess each pair’s Visual Quality, Motion Quality, Text Alignment, and Overall Quality, providing preference labels. Ultimately, VideoGen-RewardBench includes 12 T2V models, 420 prompts, and 26.5k pairs. This benchmark represents human preferences for state-of-the-art models, with videos featuring higher resolutions (480x720 - 576x1024), longer durations (4s - 6s), and improved quality. We use VideoGen-RewardBench as the primary benchmark to evaluate reward models’ performance on modern T2V models.
D.2 Comparison Methods
@@
[…]
@@
VideoScore (He et al., 2024) adopts Mantis-Idefics2-8B (Jiang et al., 2024a) as its base model and trains with point-wise data using MSE loss to model human preference scores. Since VideoScore predicts scores across multiple dimensions, and its dimension definitions differ from those in VideoGen-RewardBench, we compute both the overall accuracy and the dimension-specific (VQ, MQ, TA) accuracy by averaging the scores of five dimensions when conduct evaluation GenAI-Bench and VideoGen-RewardBench, consistent with the evaluation strategy outlined in their paper. The training data for VideoScore predominantly comes from pre-SOTA-era models, which explains its relatively better performance on GenAI-Bench, while accounts for the significant performance drop on VideoGen-RewardBench.
@@
LiFT (Wang et al., 2024b) adopts VILA-1.5-40B (Lin et al., 2024a) as its base model and employs a VLM-as-a-judge approach. The reward model is trained through instruction tuning with inputs, preference scores along with a critic. The model generates video scores and reasons through next-token prediction. LiFT evaluates videos across three dimensions: Video Fidelity, Motion Smoothness, and Semantic Consistency, which are similar to the dimensions defined in VideoGen-RewardBench. We calculate the overall accuracy using the average scores of these three dimensions and compute the dimension-specific accuracy using the corresponding dimensional scores. LiFT predicts discrete scores on a 1-3 scale, which often leads to ties in pairwise comparisons. When calculating accuracy without ties, we randomly convert the predicted tie labels to chosen/rejected with a 50% probability, indicating that the model is unable to distinguish the relative quality between the two samples.
@@
VisionReward (Xu et al., 2024a) adopts CogVLM2-Video-12B(Hong et al., 2024) as the base model and is trained to answer a set of judgment questions about the video with a binary ”yes” or ”no” response using cross-entropy loss. During inference, VisionReward evaluates 64 checklist items, providing converted into 1/0 scores. The final score is computed as the weighted average of these individual responses. We use the final score to calculate both the overall accuracy and the VQ/MQ/TA accuracy. VisionReward’s training data includes models from the pre-SOTA era models (Chen et al., 2024) as well as recent open-source T2V models (Zheng et al., 2024; Yang et al., 2024b). It performs well on GenAI-Bench and demonstrates reasonable capabilities on VideoGen-RewardBench.
@@
We adopts QWen2-VL-2B (Wang et al., 2024a) as the base model and train it with pair-wise data using BTT loss in Eq. 6. Scores are normalized on the validation set and averaged to obtain overall scores for evaluation and optimization. When evaluating on VideoGen-RewardBench, we sample videos at 2 FPS and a resolution of 448$\times$448, consistent with the training settings. We calculate the overall accuracy by averaging the scores across the three dimensions, and compute dimension-specific accuracies using the respective scores. For GenAI-Bench, we sample videos at 2 FPS and a resolution of 256$\times$256, as the minimum resolution in GenAI-Bench is 256$\times$256. Given the significant disparities in visual quality and motion between the GenAI-Bench videos and our training data, we utilize only the predicted TA scores to calculate the overall score.
D.3 Evaluation Metrics
Similarly to VisionReward (Xu et al., 2024a), we report two accuracy metrics: ties-included accuracy (Deutsch et al., 2023) and ties-excluded accuracy. For ties-excluded accuracy, we exclude all data labeled as ”ties” and use only data labeled as ”A wins” or ”B wins” for calculation. Since all competitors predict scores based on pointwise samples, we compute the rewards for each pair, convert the relative reward relationships into binary labels, and calculate classification accuracy. For ties-included accuracy, we adopt the $O(n^{2}\log n)$ algorithm proposed in Algorithm 1 by Deutsch et al. (2023). This method traverses all possible tie thresholds, calculates three-class accuracy for each threshold, and selects the highest accuracy as the final metric.
Appendix E Pseudo-code of Flow-DPO and Flow-NRG
def loss(model, ref_model, x_w, x_l, c, beta):
# ref_model: Frozen initialization of the model
# x_w: Preferred Image (latents in this work)
# x_l: Non-Preferred Image (latents in this work)
# c: Conditioning (text in this work)
# beta: Regularization Parameter
# returns: DPO loss value
timestep = torch.rand(len(x_w))
noise = torch.randn_like(x_w)
noisy_x_w = (1 - timestep) * x_w + timestep * noise
noisy_x_l = (1 - timestep) * x_l + timestep * noise
velocity_w_pred = model(noisy_x_w, c, timestep)
velocity_l_pred = model(noisy_x_l, c, timestep)
velocity_ref_w_pred = ref_model(noisy_x_w, c, timestep)
velocity_ref_l_pred = ref_model(noisy_x_l, c, timestep)
velocity_w = noise - x_w
velocity_l = noise - x_l
model_w_err = (velocity_w_pred - velocity_w).norm().pow(2)
model_l_err = (velocity_l_pred - velocity_l).norm().pow(2)
ref_w_err = (velocity_ref_w_pred - velocity_w).norm().pow(2)
ref_l_err = (velocity_ref_l_pred - velocity_l).norm().pow(2)
w_diff = model_w_err - ref_w_err
l_diff = model_l_err - ref_l_err
inside_term = -0.5 * beta * (w_diff - l_diff)
loss = -1 * log(sigmoid(inside_term))
return loss
def guidance(model, reward_model, prompt_embeds, latents, timesteps, reward_weight):
# prompt_embeds: Embeddings of the text prompts
# latents: Initial noise
# timesteps: Sequence of timesteps
# guidance_scale: scale factor for reward guidance
dts = timesteps[:-1] - timesteps[1:]
for i, t in enumerate(timesteps):
v_pred = model(latents, prompt_embeds, t)
latents = latents.detach().requires_grad_(True)
reward = reward_model(latents, prompt_embeds, t)
reward = reward * reward_weight
reward_guidance = torch.autograd.grad(reward, latents)
v_pred = v_pred - guidance_scale * t / (1 - t) * reward_guidance
latents = latents - dts[i] * v_pred
return latents
Appendix F Hyperparameters

### v2

@@ B.1 Relation beween Velocity and Noise
Let $X_{0}\sim q$ be a real data sample drawn from the true data distribution and $X_{1}\sim p$ be a noise sample, where $X_{0},X_{1}\in\mathbb{R}^{d}$. Define ${\bm{v}}_{t}({\bm{x}}_{t}\mid X_{0},X_{1})$ to be the conditional velocity field specified by a Rectified Flow Liu et al. [2022], and let ${\bm{v}}_{t}^{\text{pred}}({\bm{x}}_{t})$ be the predicted marginal velocity field. Then the L2 error of the noise prediction is related to the L2 error of the velocity field prediction by
@@
$=\;$ $\mathbb{E}\bigl[\dot{\psi}(X_{t}\mid X_{0},X_{1})\,\bigm|\,X_{t}={\bm{x}}_{t}\bigr]$
$=\;$ $\mathbb{E}\bigl[X_{1}-X_{0}\,\bigm|\,X_{t}={\bm{x}}_{t}\bigr]$
$=\;$ $\mathbb{E}\Bigl[\frac{\,X_{1}-\bigl((1-t)\,X_{0}+t\,X_{1}\bigr)\,}{\,1-t\,}\,\Bigm|\,X_{t}={\bm{x}}_{t}\Bigr]$
$=\;$ $\mathbb{E}\Bigl[\frac{\,X_{1}-{\bm{x}}_{t}\,}{\,1-t\,}\,\Bigm|\,X_{t}={\bm{x}}_{t}\Bigr]$
$=\;$ $\frac{\mathbb{E}[\,X_{1}\mid X_{t}={\bm{x}}_{t}\,]-{\bm{x}}_{t}}{\,1-t\,}.$
@@
B.2 Reward as Classifier Guidance
We begin by citing a lemma from the Guided Flows paper Zheng et al. [2023].
Lemma B.2.
@@ B.2 Reward as Classifier Guidance
Seed Appendix A.61 of the Guided Flows paper Zheng et al. [2023] for detailed derivations.
@@
$\tilde{v}_{t}({\bm{x}}_{t}|{\bm{y}})$ $={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[a_{t}{\bm{x}}_{t}+b_{t}\nabla r({\bm{x}}_{t},{\bm{y}})-{\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})]$
$={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[a_{t}{\bm{x}}_{t}+b_{t}\nabla\log\exp(r({\bm{x}}_{t},{\bm{y}}))-{\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})]$
$=(1-w){\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[a_{t}{\bm{x}}_{t}+b_{t}\nabla\log\exp(r({\bm{x}}_{t},{\bm{y}}))]$
$=a_{t}{\bm{x}}_{t}+b_{t}\nabla\left[(1-w)\log p_{t}({\bm{x}}_{t}|{\bm{y}})+w\log\exp(r({\bm{x}}_{t},{\bm{y}}))\right]$
@@
We note that this analysis shows that Reward Guided Flows are guaranteed to sample from $\tilde{q}(\cdot|{\bm{y}})$ at time $t=1$ if the probability path $\tilde{p}_{t}(\cdot|{\bm{y}})$ is close to the marginal probability path $\int p_{t}(\cdot|{\bm{x}}_{1})\tilde{q}({\bm{x}}_{1}|{\bm{y}})dx_{1}$, but it is not clear to what extent this assumption holds in practice. This also mens that $\tilde{p_{t}}({\bm{x}}_{t}|{\bm{y}})$ is also a marginal gaussian path defined by $p_{t}({\bm{x}}|{\bm{x}}_{1})=\mathcal{N}({\bm{x}}|\alpha_{t}{\bm{x}}_{1},\sigma_{t}^{2}I)$.
@@
$\tilde{v}_{t}({\bm{x}}_{t}|{\bm{y}})$ $={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+wb_{t}\nabla r({\bm{x}}_{t},{\bm{y}})$
$={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+wb_{t}\nabla\log\exp(r({\bm{x}}_{t},{\bm{y}}))$
$=a_{t}{\bm{x}}_{t}+b_{t}\nabla\left[\log p_{t}({\bm{x}}_{t}|{\bm{y}})+w\log\exp(r({\bm{x}}_{t},{\bm{y}}))\right]$
@@
Rectified Flow Liu et al. [2022] is also a Gaussian path defined by
@@
We use Eq. 17 or Eq. B.2 to guide inference through reward model.
Appendix C Pseudo-code of Flow-DPO and Flow-NRG
def loss(model, ref_model, x_w, x_l, c, beta):
# ref_model: Frozen initialization of the model
# x_w: Preferred Image (latents in this work)
# x_l: Non-Preferred Image (latents in this work)
# c: Conditioning (text in this work)
# beta: Regularization Parameter
timestep = torch.rand(len(x_w))
noise = torch.randn_like(x_w)
noisy_x_w = (1 - timestep) * x_w + timestep * noise
noisy_x_l = (1 - timestep) * x_l + timestep * noise
velocity_w_pred = model(noisy_x_w, c, timestep)
velocity_l_pred = model(noisy_x_l, c, timestep)
velocity_ref_w_pred = ref_model(noisy_x_w, c, timestep)
velocity_ref_l_pred = ref_model(noisy_x_l, c, timestep)
velocity_w = noise - x_w
velocity_l = noise - x_l
model_w_err = (velocity_w_pred - velocity_w).norm().pow(2)
model_l_err = (velocity_l_pred - velocity_l).norm().pow(2)
ref_w_err = (velocity_ref_w_pred - velocity_w).norm().pow(2)
ref_l_err = (velocity_ref_l_pred - velocity_l).norm().pow(2)
w_diff = model_w_err - ref_w_err
l_diff = model_l_err - ref_l_err
inside_term = -0.5 * beta * (w_diff - l_diff)
loss = -1 * log(sigmoid(inside_term))
return loss
def reward_guidance(model, reward_model, prompt_embeds, latents, timesteps, reward_weight, rg_scale, cfg_scale):
# prompt_embeds: Embeddings of the text prompts
# latents: Initial noise
# timesteps: Sequence of timesteps
# rg_scale: scale factor for reward guidance
# cfg_scale: scale factor for classifier free guidance
dts = timesteps[:-1] - timesteps[1:]
for i, t in enumerate(timesteps):
v_pred = model(latents, prompt_embeds, t)
if cfg_scale != 1.0:
v_pred_uncond = model(latents, None, t) # unconditional prediction
v_pred = v_pred_uncond + cfg_scale * (v_pred - v_pred_uncond)
latents = latents.detach().requires_grad_(True)
reward = reward_model(latents, prompt_embeds, t)
reward = (reward * reward_weight).sum()
reward_guidance = torch.autograd.grad(reward, latents)
if t != 1:
v_pred = v_pred - rg_scale * t / (1 - t) * reward_guidance
latents = latents - dts[i] * v_pred
return latents
Appendix D Architecture of Internal Video Diffusion Model
Our work employs an internal text-to-video foundation model, which is a Transformer-based latent diffusion model Peebles and Xie [2023] with 1B parameters. The model integrates a 3D VAE to encode video data into a latent space, alongside a Transformer-based video diffusion model. Each Transformer block is instantiated as a sequence of 2D spatial-attention, 3D self-attention, cross-attention, and FFN modules. Importantly, the model is trained under Flow Matching framework. An illustration of the model architecture is provided in Fig 7.
Figure 7: Architecture of Our Internal Video Latent Diffusion Model Backbone.
Appendix E Ablation Study of Our Reward Model
Ablation of VideoReward
To further validate the key design choices of our reward model discussed in Sec.3.2, we conduct an ablation study on the evaluation benchmark, providing a quantitative supplement to our analysis in Tab. 6. We compare three reward model variants: regression-style, Bradley-Terry, and Bradley-Terry-With-Tied. The BT model slightly outperforms the regression-style model, likely due to the advantages of pairwise annotations, which better capture ranking relationships and are more robust to annotation noise. The BTT model matches the BT model on ties-excluded accuracy but significantly improves ties-included accuracy, as its explicit handling of tied pairs helps it learn a more robust decision boundary, capturing neutral relationships in ambiguous cases. Additionally, we find that using separate tokens for each reward attribute, instead of a shared last token further improves performance. This design better represents distinct reward aspects, enhancing alignment with human preferences.
Table 6: Ablation study on reward model type and seprate tokens. Bold: Best Performance.
Variants RM Type Separate Tokens GenAI-Bench VideoGen-RewardBench
II BT 47.74 71.21 61.22 72.58 52.33 77.10 59.43 73.50 53.06 71.62
III BTT 48.27 70.89 61.50 73.39 60.52 76.31 64.64 72.40 53.55 72.12
IV BTT 49.41 72.89 61.26 73.59 59.68 75.66 66.03 74.70 53.80 72.20
Ablation on Noisy-Latent Reward Training.
We apply reward guidance using only MQ rewards on TA-Hard prompts. When trained with noised latents, the reward model successfully guides generation, yielding an MQ win rate of 74.6. By contrast, the model trained on clean latents fails to offer meaningful guidance for intermediate noised latents, achieving only a 38.6 win rate and underperforming even the unguided baseline. Following Yu et al. [2023], we also try directly predicting ${\bm{x}}_{0}$ from ${\bm{x}}_{t}$ and then backpropagating via $r({\bm{x}}_{0},y)$. However, this method produces unusable videos without meaningful content.
Appendix F Statistics of Training Data
In Tab. 1 and Fig. 8, we provide a comprehensive statistics of our 182k-sized training dataset.
Figure 8: Statistics of our training data.
Previous reward models (e.g., VideoScore He et al. [2024], LIFT Wang et al. [2024b]) were trained on datasets with mostly low resolutions ($\leq$480p) and short durations (2s), whereas ours include recent T2V models with higher resolutions ($\geq$720p) and longer clips (3–6s). The dataset statistics are shown in Tab. 7.
Table 7: Comparison of video preference datasets in terms of spatial resolution and temporal duration.
Dataset Resolution Duration
Existing Preference Datasets
VideoFeedback He et al. [2024] $256\times 256$$576\times 1024$ 1s – 3s
VideoDPO Liu et al. [2024b] $320\times 512$ 2s
LiFT-HRA Wang et al. [2024b] $480\times 720$ 6s
Our Datasets
Ours $384\times 672$$768\times 1408$ 3s – 6s
Appendix G Details of Human Annotation
@@
Appendix H Details of Reward Model Evaluation
H.1 Evaluation Benchmarks
We evaluate our reward model using two benchmarks: GenAI-Bench Jiang et al. [2024b] and VideoGen-RewardBench. GenAI-Bench is employed to assess the accuracy of the reward model in evaluating pre-SOTA-era T2V models, while VideoGen-RewardBench is used to evaluate its performance on modern T2V models. In this subsection, we describe both benchmarks, highlighting key parameters and differences in Fig 9 and Tab. 9. We also visualize the model coverage across the training sets of different baselines and the two evaluation benchmarks, as shown in the Fig 10.
@@ H.1 Evaluation Benchmarks
VideoGen-Eval Zeng et al. [2024] has open-sourced a dataset containing videos generated by 9 closed-source and 3 open-source models, designed to qualitatively visualize performance differences across models. Due to its high-quality data, broad coverage of the latest advancements in T2V models, and third-party nature, we leverage VideoGen-Eval to create a fair benchmark, VideoGen-RewardBench, for evaluating reward models’ performance on modern T2V models. We manually construct 26.5k video pairs and hire annotators to assess each pair’s Visual Quality, Motion Quality, Text Alignment, and Overall Quality, providing preference labels. Ultimately, VideoGen-RewardBench includes 12 T2V models, 420 prompts, and 26.5k pairs. This benchmark represents human preferences for state-of-the-art models, with videos featuring higher resolutions (480x720 - 576x1024), longer durations (4s - 6s), and improved quality. We use VideoGen-RewardBench as the primary benchmark to evaluate reward models’ performance on modern T2V models.
H.2 Comparison Methods
@@
To eliminate the influence of metric calculations and benchmark distributions on our evaluation results, we introduce a special baseline: random scores. Specifically, for each triplet (prompt, video A, video B), we randomly sample $r_{A}$ and $r_{B}$ from a standard normal distribution, denoted as $r_{A},r_{B}\sim\mathcal{N}(0,1)$. We then calculate accuracy in the same manner as for the other models. The mathematical expectation of random scores for ties-excluded accuracy is $\mathbb{E}(acc)=\frac{1}{2}$, and the mathematical expectation of ties-included accuracy is $\mathbb{E}(acc)=\max(\frac{1}{3},p(c=\text{"Ties"}))$.
@@
VideoScore He et al. [2024] adopts Mantis-Idefics2-8B Jiang et al. [2024a] as its base model and trains with point-wise data using MSE loss to model human preference scores. Since VideoScore predicts scores across multiple dimensions, and its dimension definitions differ from those in VideoGen-RewardBench, we compute both the overall accuracy and the dimension-specific (VQ, MQ, TA) accuracy by averaging the scores of five dimensions when conduct evaluation GenAI-Bench and VideoGen-RewardBench, consistent with the evaluation strategy outlined in their paper. The training data for VideoScore predominantly comes from pre-SOTA-era models, which explains its relatively better performance on GenAI-Bench, while accounts for the significant performance drop on VideoGen-RewardBench.
@@
LiFT Wang et al. [2024b] adopts VILA-1.5-40B Lin et al. [2024a] as its base model and employs a VLM-as-a-judge approach. The reward model is trained through instruction tuning with inputs, preference scores along with a critic. The model generates video scores and reasons through next-token prediction. LiFT evaluates videos across three dimensions: Video Fidelity, Motion Smoothness, and Semantic Consistency, which are similar to the dimensions defined in VideoGen-RewardBench. We calculate the overall accuracy using the average scores of these three dimensions and compute the dimension-specific accuracy using the corresponding dimensional scores. LiFT predicts discrete scores on a 1-3 scale, which often leads to ties in pairwise comparisons. When calculating accuracy without ties, we randomly convert the predicted tie labels to chosen/rejected with a 50% probability, indicating that the model is unable to distinguish the relative quality between the two samples.
@@
VisionReward Xu et al. [2024a] adopts CogVLM2-Video-12BHong et al. [2024] as the base model and is trained to answer a set of judgment questions about the video with a binary ”yes” or ”no” response using cross-entropy loss. During inference, VisionReward evaluates 64 checklist items, providing converted into 1/0 scores. The final score is computed as the weighted average of these individual responses. We use the final score to calculate both the overall accuracy and the VQ/MQ/TA accuracy. VisionReward’s training data includes models from the pre-SOTA era models Chen et al. [2024b] as well as recent open-source T2V models Zheng et al. [2024], Yang et al. [2024b]. It performs well on GenAI-Bench and demonstrates reasonable capabilities on VideoGen-RewardBench.
@@
We adopts QWen2-VL-2B Wang et al. [2024a] as the base model and train it with pair-wise data using BTT loss in Eq. 2. Scores are normalized on the validation set and averaged to obtain overall scores for evaluation and optimization. When evaluating on VideoGen-RewardBench, we sample videos at 2 FPS and a resolution of $\times448$, consistent with the training settings. We calculate the overall accuracy by averaging the scores across the three dimensions, and compute dimension-specific accuracies using the respective scores. For GenAI-Bench, we sample videos at 2 FPS and a resolution of $\times256$, as the minimum resolution in GenAI-Bench is $\times256$. Given the significant disparities in visual quality and motion between the GenAI-Bench videos and our training data, we utilize only the predicted TA scores to calculate the overall score.
H.3 Evaluation Metrics
Similarly to VisionReward Xu et al. [2024a], we report two accuracy metrics: ties-included accuracy Deutsch et al. [2023] and ties-excluded accuracy. For ties-excluded accuracy, we exclude all data labeled as ”ties” and use only data labeled as ”A wins” or ”B wins” for calculation. Since all competitors predict scores based on pointwise samples, we compute the rewards for each pair, convert the relative reward relationships into binary labels, and calculate classification accuracy. For ties-included accuracy, we adopt the tie calibration algorithm proposed in Algorithm 1 by Deutsch et al. [2023]. This method traverses all possible tie thresholds, calculates three-class accuracy for each threshold, and selects the highest accuracy as the final metric.
Appendix I Hyperparameters
In all alignment experiments, we applied LoRA to fine-tune the transformer models’ linear layers, as our findings indicate that full parameter fine-tuning can degrade the model’s performance or potentially lead to model collapse.
Table 10: Hyperparameters for alignment algorithms